# A unified convention for achievement positional games

### v1

@@ 3.1 Disjoint union
If $\mathcal{G}=(V,E_{L},E_{R})$ and $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ are two achievement positional games such that $V$ and $V^{\prime}$ are disjoint, we define their disjoint union $\mathcal{G}\cup\mathcal{G}^{\prime}=(V\cup V^{\prime},E_{L}\cup E^{\prime}_{L},E_{R}\cup E^{\prime}_{R})$. As with all conventions of positional games, a natural question is that of the outcome of a disjoint union given the outcomes of both components. This is one of the reasons why, similarly to what is done in combinatorial game theory [BCG82], we have opted for a definition of the outcome which does not assume that, say, Left always starts. Indeed, when considering a disjoint union, both cases are relevant since even the first player will basically act as second player on whichever component they choose not to start in.
In the Maker-Maker convention, it is known that the disjoint union of two hypergraphs that are first player wins is also a first player win. This fact is not straightforward, as the second player might not compliantly play each move in the same component the first player has just played in. However, there exists a simple proof based on comparing the delay of both components, which is an adequate measure of how fast the first player can win [Lea08]. We now generalize this argument to achievement positional games.
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game such that Left has a winning strategy on $\mathcal{G}$ as first player. We define the Left-delay of $\mathcal{G}$ through the following scoring game denoted by $sc_{L}(\mathcal{G})$. Consider the game $\mathcal{G}$ except that Right plays second and has the option, on every turn, to either play normally or pass her move. Left plays first and acts in the usual way. As normal, the game ends when Left fills a blue edge, or Right fills a red edge, or all vertices have been picked. The score of the game is +$\infty$+ ∞ if Left has not filled a blue edge, or is equal to the total number of pass moves made by Right otherwise. The Left-delay of $\mathcal{G}$ is defined as the score of the game when both players play optimally, meaning that Left aims at minimizing the score while Right aims at maximizing it. Note that the Left-delay of $\mathcal{G}$ is always finite since we assume that Left has a winning strategy on $\mathcal{G}$ as first player. If Right has a winning strategy on $\mathcal{G}$ as first player, then we define the scoring game $sc_{R}(\mathcal{G})$ and the Right-delay of $\mathcal{G}$ in an analogous manner.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ and $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ be two achievement positional games with disjoint vertex sets. We suppose that Left has a winning strategy on $\mathcal{G}$ as first player, and that Right has a winning strategy on $\mathcal{G}^{\prime}$ as first player. Then either Left or Right (or both) has a winning strategy as first player on $\mathcal{G}\cup\mathcal{G}^{\prime}$. More precisely, denoting by $d$ the Left-delay of $\mathcal{G}$ and by $d^{\prime}$ the Right-delay of $\mathcal{G}^{\prime}$:
  • If $d\leq d^{\prime}$, then Left has a winning strategy on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as first player.
  • If $d^{\prime}\leq d$, then Right has a winning strategy on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as first player.
@@
We now define Left’s winning strategy $\widetilde{\mathcal{S}}$ as first player on $\mathcal{G}\cup\mathcal{G}^{\prime}$. He starts by playing in the component $\mathcal{G}$, according to $\mathcal{S}$. After this, Left always follows $\mathcal{S}$ or $\mathcal{S}^{\prime}$, depending on which component Right has just played in. We do need to address two situations that may occur when Right plays in $\mathcal{G}^{\prime}$:
@@
[…]
@@
We start by showing that, for each cell of Table 3, all outcomes that are not listed are indeed impossible. Recall that, by Lemma 2.2, playing twice in a row in the same component cannot harm any player’s chances in that component. Therefore, it will often be implicitly assumed that, if a player was supposed to answer in the same component as their opponent but cannot do so because there is no vertex left in that component, then that player can pick an arbitrary vertex of the other component instead. Using symmetries, proving the following claims is sufficient.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as second player. Left can simply play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}^{\prime}$, and will win in $\mathcal{G}$.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as first player. Left can simply start in $\mathcal{G}^{\prime}$ and then play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}^{\prime}$, and will win in $\mathcal{G}$.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as first player. Left can simply start in $\mathcal{G}^{\prime}$ and then play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}$, and will win in $\mathcal{G}^{\prime}$.
@@
    The proof that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as first player is the same as the previous case. Moreover, as second player, Left can at least draw the game by playing each move in the same component Right has just played in, according to his non-losing strategies in each component.
@@
    The fact that $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{D}$ is a direct consequence of Proposition 3.1. Moreover, as first player, Left can at least draw the game by starting in $\mathcal{G}^{\prime}$ and then playing each move in the same component Right has just played in, according to his non-losing strategies in each component. This proves that $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{R}$. Similarly, $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{L}$.
@@ 3.2 Elementary strategic principles
The goal of this subsection is to list simple general properties of achievement positional games. Even though some nice properties of the Maker-Breaker convention disappear when considering the Maker-Maker convention instead, we are about to see that not much more seems to be lost when generalizing to achievement positional games.
A well-known strategy-stealing argument [HJ63] ensures that the second player can never have a winning strategy for the Maker-Maker convention. As we know, when it comes to achievement positional games in general, unbalance between $E_{L}$ and $E_{R}$ can lead to a one-sided outcome $\mathcal{L}$ or $\mathcal{R}$. However, if $E_{L}$ is “at least as good” as $E_{R}$ in some adequate sense for instance, then Left should have a non-losing strategy as first player. We show that this is indeed the case. Before that, let us remark that $(V,E_{L})$ and $(V,E_{R})$ being isomorphic is not a sufficient condition, as Figure 4 features a counterexample with outcome $\mathcal{R}$. Indeed, Right’s red edges of size 2 allow her to destroy the blue butterflies, and one of her own red butterflies will be left intact for her to win the game in the end.
Figure 4: Despite the blue and red hypergraphs being isomorphic, Right has a winning strategy as second player.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. If there exists a bijection $\sigma:V\to V$ such that $\sigma(e)\in E_{L}$ and $\sigma^{-1}(e)\in E_{L}$ for all $e\in E_{R}$, then Left has a non-losing strategy on $\mathcal{G}$ as first player.
@@
Suppose for a contradiction that Right has a winning strategy $\mathcal{S}$ on $\mathcal{G}$ as second player. We interpret $\sigma$ as a mapping from a fictitious game, where Left starts and loses against the strategy $\mathcal{S}$, to the real game, where Left starts and wins using a strategy $\mathcal{S}^{\prime}$ which we now describe. Left starts by picking an arbitrary vertex $a_{1}$ in the real game, and then acts in the following way for all $i\geq 1$ until the game ends:
@@
  • In the fictitious game, Left picks $\alpha_{i}=\sigma^{-1}(b_{i})$, to which Right has a winning answer $\beta_{i}$ according to $\Sigma$.
@@
All in all, $\mathcal{S}^{\prime}$ is a winning strategy for Left on $\mathcal{G}$ as first player, contradicting the existence of $\mathcal{S}$. ∎
Consider Hex for example: Lemma 3.3 applied to the diagonal mirror symmetry $\sigma$ ensures that the first player has a non-losing strategy (see Figure 5). As it is technically impossible for the game to end in a draw [Gar59], this implies that the first player has a winning strategy.
@@
We now address monotonicity. A key feature of the Maker-Breaker convention is that adding more edges cannot harm Maker. One of the main reasons why the Maker-Maker convention is harder to study is because, in contrast, adding an edge might transform a first player win into a draw: this is known as the extra set paradox [Bec08]. However, this phenomenon is due to the fact that the extra edge is there for both players to fill. In our general context of achievement positional games, adding or shrinking blue edges cannot harm Left, and adding or shrinking red edges cannot harm Right.
@@
Pairing strategies are an important tool in both Maker-Breaker and Maker-Maker conventions [HJ63]. A complete pairing of a hypergraph $H$ is a set $\Pi$ of pairwise disjoint pairs of vertices such that every edge of $H$ contains some element of $\Pi$. If $H$ admits a complete pairing, then the outcome is a Breaker win for the Maker-Breaker game or a draw for the Maker-Maker game, as picking one vertex from each pair prevents the other player from filling an edge. We observe that, in general achievement positional games, pairing strategies may still be used to block the opponent.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. If the hypergraph $(V,E_{L})$ admits a complete pairing $\Pi$, then Right has a non-losing strategy on $\mathcal{G}$ as first or second player.
@@
It should be noted that, in the Maker-Breaker convention, the same conclusion holds under the weaker assumption that, for all $S\subseteq V\setminus\{u,v\}$, we have $S\cup\{u\}\in E\iff S\cup\{v\}\in E$ (see [Oij24, Lemma 1.84]). For instance, this allows to simplify two edges $\{u,v,w_{1}\}$ and $\{u,v,w_{2}\}$ into a single edge $\{u,v\}$ if $w_{1}$ and $w_{2}$ are of degree 1. This is already false in the Maker-Maker convention, where the notion of speed becomes important. All in all, even though Lemma 3.7 admits a stronger version for the Maker-Breaker convention, it likely does not for the Maker-Maker convention.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. Suppose that there is no edge of size 1, but there is a blue (resp. red) edge $\{u,v\}$ such that, for all $e\in E_{L}\cup E_{R}:u\in e\implies v\in e$. Then it is optimal for Left (resp. Right) as first player to start by picking $v$, which forces the opponent to answer by picking $u$.
@@
Let us address Left’s point of view: the updated game after this first round of play is $\mathcal{G}_{v}^{u}$, and we have $o(\mathcal{G})\leq_{L}o(\mathcal{G}_{v}^{u})$ by item (iii) from Lemma 3.6. ∎

### v2

@@ 3.1 Disjoint union
If $\mathcal{G}=(V,E_{L},E_{R})$ and $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ are two achievement positional games such that $V$ and $V^{\prime}$ are disjoint, we define their disjoint union $\mathcal{G}\cup\mathcal{G}^{\prime}=(V\cup V^{\prime},E_{L}\cup E^{\prime}_{L},E_{R}\cup E^{\prime}_{R})$. As with all conventions of positional games, a natural question is that of the outcome of a disjoint union given the outcomes of both components. This is one of the reasons why, similarly to what is done in combinatorial game theory [3], we have opted for a definition of the outcome which does not assume that, say, Left always starts. Indeed, when considering a disjoint union, both cases are relevant since even the first player will basically act as the second player on whichever component they choose not to start in.
In the Maker-Maker convention, it is known that the disjoint union of two hypergraphs that are first player wins is also a first player win. This fact is not straightforward, as the second player might not compliantly play each move in the same component the first player has just played in. However, there exists a simple but little-known proof based on comparing the delay of both components, which is an adequate measure of how fast the first player can win [16]. We now generalize this argument to achievement positional games.
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game such that Left has a winning strategy on $\mathcal{G}$ as the first player. We define the Left-delay of $\mathcal{G}$ through the following scoring game denoted by $sc_{L}(\mathcal{G})$. Consider the game $\mathcal{G}$ except that Right plays second and has the option, on every turn, to either play normally or pass her move. Left plays first and acts in the usual way. As normal, the game ends when Left fills a blue edge, or Right fills a red edge, or all vertices have been picked. The score of the game is $\infty$ if Left has not filled a blue edge, or is equal to the total number of pass moves made by Right otherwise. The Left-delay of $\mathcal{G}$ is defined as the score of the game when both players play optimally, meaning that Left aims at minimizing the score while Right aims at maximizing it. Note that the Left-delay of $\mathcal{G}$ is always finite since we assume that Left has a winning strategy on $\mathcal{G}$ as the first player. If Right has a winning strategy on $\mathcal{G}$ as the first player, then we define the scoring game $sc_{R}(\mathcal{G})$ and the Right-delay of $\mathcal{G}$ in an analogous manner.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ and $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ be two achievement positional games with disjoint vertex sets. We suppose that Left has a winning strategy on $\mathcal{G}$ as the first player, and that Right has a winning strategy on $\mathcal{G}^{\prime}$ as the first player. Then either Left or Right (or both) has a winning strategy as the first player on $\mathcal{G}\cup\mathcal{G}^{\prime}$. More precisely, denoting by $d$ the Left-delay of $\mathcal{G}$ and by $d^{\prime}$ the Right-delay of $\mathcal{G}^{\prime}$:
  • If $d\leq d^{\prime}$, then Left has a winning strategy on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the first player.
  • If $d^{\prime}\leq d$, then Right has a winning strategy on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the first player.
@@
We now define Left’s winning strategy $\widetilde{\mathcal{S}}$ as the first player on $\mathcal{G}\cup\mathcal{G}^{\prime}$. He starts by playing in the component $\mathcal{G}$, according to $\mathcal{S}$. After this, Left always obeys $\mathcal{S}$ or $\mathcal{S}^{\prime}$, depending on which component Right has just played in. We do need to address two situations that may occur when Right plays in $\mathcal{G}^{\prime}$:
@@
$\mathcal{G}^{\prime}$ $\mathcal{G}$ $\mathcal{L}$ $\mathcal{L}^{-}$ $\mathcal{N}$ $\mathcal{D}$ $\mathcal{R}^{-}$ $\mathcal{R}$
$\mathcal{L}$ $\mathcal{L}$ $\mathcal{L}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$
$\mathcal{L}^{-}$ $\mathcal{L}$ $\mathcal{L}$, $\mathcal{L}^{-}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}^{-}$ $\mathcal{L}^{-}$, $\mathcal{N}$, $\mathcal{R}^{-}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$
$\mathcal{N}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{N}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$
$\mathcal{D}$ $\mathcal{L}$ $\mathcal{L}^{-}$ $\mathcal{N}$ $\mathcal{D}$ $\mathcal{R}^{-}$ $\mathcal{R}$
$\mathcal{R}^{-}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$ $\mathcal{L}^{-}$, $\mathcal{N}$, $\mathcal{R}^{-}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{R}^{-}$ $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{R}$
$\mathcal{R}$ $\mathcal{L}$, $\mathcal{L}^{-}$, $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{N}$, $\mathcal{R}^{-}$, $\mathcal{R}$ $\mathcal{R}$ $\mathcal{R}$ $\mathcal{R}$
@@
We start by showing that, for each cell of Table 3, all outcomes that are not listed are indeed impossible. Recall that, by Lemma 2.3, playing twice in a row in the same component cannot harm any player’s chances in that component. Therefore, it will often be implicitly assumed that, if a player was supposed to answer in the same component as their opponent but cannot do so because there is no vertex left in that component, then that player picks an arbitrary vertex of the other component instead. Using symmetries, proving the following claims is sufficient.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the second player. Left can simply play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}^{\prime}$, and will win in $\mathcal{G}$.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the first player. Left can simply start in $\mathcal{G}^{\prime}$ and then play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}^{\prime}$, and will win in $\mathcal{G}$.
@@
    To prove this, we must show that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the first player. Left can simply start in $\mathcal{G}^{\prime}$ and then play each move in the same component Right has just played in, according to his optimal strategies in each component. Doing so, Left will not lose in $\mathcal{G}$, and will win in $\mathcal{G}^{\prime}$.
@@
    The proof that Left wins on $\mathcal{G}\cup\mathcal{G}^{\prime}$ as the first player is the same as the previous case. Moreover, as the second player, Left can at least draw the game by playing each move in the same component Right has just played in, according to his non-losing strategies in each component.
@@
    The fact that $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{D}$ is a direct consequence of Proposition 3.1. Moreover, as the first player, Left can at least draw the game by starting in $\mathcal{G}^{\prime}$ and then playing each move in the same component Right has just played in, according to his non-losing strategies in each component. This proves that $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{R}$. Similarly, $o(\mathcal{G}\cup\mathcal{G}^{\prime})\neq\mathcal{L}$.
@@ 3.2 Elementary strategic principles
The goal of this subsection is to list simple general properties of achievement positional games. Even though some nice properties of the Maker-Breaker convention disappear when considering the Maker-Maker convention instead, we are about to see that not much more seems to be lost when generalizing to achievement positional games. Each result is only stated from one player’s perspective, but the mirror result when swapping Left with Right and $E_{L}$ with $E_{R}$ obviously holds as well.
A well-known strategy-stealing argument [13] ensures that the second player can never have a winning strategy for the Maker-Maker convention. As we know, when it comes to achievement positional games in general, unbalance between $E_{L}$ and $E_{R}$ can lead to a one-sided outcome $\mathcal{L}$ or $\mathcal{R}$. However, if $E_{L}$ is “at least as good” as $E_{R}$ in some adequate sense for instance, then Left should have a non-losing strategy as the first player. We show that this is indeed the case. Before that, let us remark that $(V,E_{L})$ and $(V,E_{R})$ being isomorphic is not a sufficient condition, as Figure 4 features a counterexample with outcome $\mathcal{R}$. Indeed, the red 2-edges allow Right to destroy the blue butterflies, and one of her own red butterflies will be left intact for her to win the game in the end.
Figure 4: Despite the blue and red hypergraphs being isomorphic, Right has a winning strategy as the second player.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. If there exists a bijection $\sigma:V\to V$ such that $\sigma(e)\in E_{L}$ and $\sigma^{-1}(e)\in E_{L}$ for all $e\in E_{R}$, then Left has a non-losing strategy on $\mathcal{G}$ as the first player.
@@
Suppose for a contradiction that Right has a winning strategy $\mathcal{S}$ on $\mathcal{G}$ as the second player. We interpret $\sigma$ as a mapping from a fictitious game, where Left starts and loses against the strategy $\mathcal{S}$, to the real game, where Left starts and wins using a strategy $\mathcal{S}^{\prime}$ which we now describe. Left starts by picking an arbitrary vertex $a_{1}$ in the real game, and then acts in the following way for all $i\geq 1$ until the game ends:
@@
  • In the fictitious game, Left picks $\alpha_{i}=\sigma^{-1}(b_{i})$, to which Right has a winning answer $\beta_{i}$ according to $\mathcal{S}$.
@@
All in all, $\mathcal{S}^{\prime}$ is a winning strategy for Left on $\mathcal{G}$ as the first player, contradicting the existence of $\mathcal{S}$. ∎
Consider Hex for example: Lemma 3.3 applied to the diagonal mirror symmetry $\sigma$ ensures that the first player has a non-losing strategy (see Figure 5). As it is technically impossible for the game to end in a draw [12], this implies that the first player has a winning strategy.
@@
We now address monotonicity. A key feature of the Maker-Breaker convention is that adding more edges cannot harm Maker. One of the main reasons why the Maker-Maker convention is harder to study is because, in contrast, adding an edge might transform a first player win into a draw: this is known as the extra set paradox [2]. However, this phenomenon is due to the fact that the extra edge is there for both players to fill. In our general context of achievement positional games, adding or shrinking blue edges cannot harm Left, and adding or shrinking red edges cannot harm Right.
@@
Pairing strategies are an important tool in both Maker-Breaker and Maker-Maker conventions [13]. A complete pairing of a hypergraph $H=(V,E)$ is a set $\Pi$ of pairwise-disjoint pairs of vertices such that, for all $e\in E$, there exists $\pi\in\Pi$ satisfying $\pi\subseteq e$. If $H$ admits a complete pairing, then the outcome of $H$ is a Breaker win for the Maker-Breaker game or a draw for the Maker-Maker game, as picking one vertex from each pair prevents the other player from filling an edge. We observe that, in general achievement positional games, pairing strategies may still be used to block the opponent.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. If the hypergraph $(V,E_{L})$ admits a complete pairing $\Pi$, then Right has a non-losing strategy on $\mathcal{G}$ as the first player and as the second player.
@@
It should be noted that, in the Maker-Breaker convention, the same conclusion holds under the weaker assumption that, for all $S\subseteq V\setminus\{u,v\}$, we have $S\cup\{u\}\in E\iff S\cup\{v\}\in E$ (see [17, Lemma 1.84]). For instance, this allows to simplify two edges $\{u,v,w_{1}\}$ and $\{u,v,w_{2}\}$ into a single edge $\{u,v\}$ if $w_{1}$ and $w_{2}$ are of degree 1. This is already false in the Maker-Maker convention, where the notion of speed becomes important. All in all, even though Lemma 3.7 admits a stronger version for the Maker-Breaker convention, it likely does not for the Maker-Maker convention.
@@
Let $\mathcal{G}=(V,E_{L},E_{R})$ be an achievement positional game. Suppose that there is no 1-edge, but there is a blue edge $\{u,v\}$ such that, for all $e\in E_{L}\cup E_{R}:u\in e\implies v\in e$. Then it is optimal for Left as the first player to start by picking $v$, which forces Right to answer by picking $u$.
@@
The updated game after this first round of play is $\mathcal{G}_{v}^{u}$, and we have $o(\mathcal{G})\leq_{L}o(\mathcal{G}_{v}^{u})$ by item (iii) from Lemma 3.6. ∎